# Oracle-Based Multistep Strategy for Solving Polynomial Systems Over Finite Fields and Algebraic Cryptanalysis of the Aradi Cipher

### v1

@@ Abstract.
In this paper, we present a new implementation of the corresponding algorithm based on a Depth-First Search strategy, along with a novel complexity analysis leveraging tree structures. We further introduce the notion of an “oracle function” as a general predictive tool for deciding whether the evaluation of a new variable is necessary to simplify the current polynomial system. This notion allows us to unify all previously proposed variants of the multistep strategy, including the classical hybrid approach, by appropriately selecting the oracle function.
Finally, we apply the multistep solving strategy to the cryptanalysis of the low-latency block cipher Aradi, recently introduced by the NSA. We present the first full-round algebraic attack, raising concerns about the cipher’s actual security with respect to its key length.
@@
Primary 13P15. Secondary 11T71, 12H10
The first author acknowledges the partial support of PNRR MUR projects “Security and Rights in the CyberSpace”, Grant ref. CUP H93C22000620001, Code PE00000014, Spoke 3, and “National Center for HPC, Big Data and Quantum Computing”, Grant ref. CUP H93C22000450007, Code CN00000013, Spoke 10. The same author was co-funded by Università di Bari “Fondo acquisto e manutenzione attrezzature per la ricerca”, Grant ref. DR 3191.
@@ 1. Introduction
Beyond theoretical worst-case complexity, valid when the polynomials in the system are randomly chosen, it is especially important in cryptanalysis to estimate the actual computational cost of solving a specific polynomial system over a finite field through experimentation. Since solving such systems becomes infeasible when the number of variables is sufficiently large, a standard approach is to combine brute force with polynomial solving in what is known as a “hybrid strategy”. Specifically, this approach assumes that one can feasibly solve all derived systems obtained by exhaustively assigning values in the finite field to $k$ out of the $n$ variables in the original system. By using a test set of such partial evaluations, it is possible to estimate the average solving time $\tau$ for a single derived system. Consequently, the total expected complexity of the hybrid approach becomes $q^{n-k}\tau$.
@@
Moreover, we observe that all derived systems are quadratic and can be solved with a solving degree of two when more than 250 key variable assignments are provided. However, this is no longer possible when the number of assignments is reduced to 250. These findings reveal thresholds within Aradi’s algebraic structure that are relevant to its security, and demonstrate the practical benefits of employing a dynamic multistep strategy for algebraic cryptanalysis.
@@ 2. Multistep algorithm
Let ${\mathbb{F}}={\mathrm{GF}}(q)$ be a finite field and consider the multivariate polynomial algebra $R={\mathbb{F}}[x_{1},\ldots,x_{n}]$ and the field equation ideal ${\mathcal{F}}=\langle x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\rangle\subset R$. Let $J=\langle f_{1},\ldots,f_{m}\rangle\subset R$ be an ideal such that #$\#V_{\mathbb{F}}(J)\leq 1$. Put
@@
that is, $H$ is a generating set of the ideal $J+{\mathcal{F}}$. To simplify notations, we denote ${\mathbb{F}}=\{c_{1},\ldots,c_{q}\}$. Since #$\#V_{\mathbb{F}}(J)\leq 1$, we have that a (reduced) Gröbner basis of $J+{\mathcal{F}}$ with respect to any monomial ordering of $R$, is either $G=\{1\}$ or $G=\{x_{1}-c_{i_{1}},\ldots,x_{n}-c_{i_{n}}\}$, for some $1\leq i_{1},\ldots,i_{n}\leq q$ (see, for instance, [14, 12]). In other words, ${\mathrm{maxdeg}}(G)=\max\{\deg(f)\mid f\in G\}\leq 1$.
Let $d_{1}\geq 0$ be an integer and denote by ${\textsc{Grobner}}(H,d_{1})$ the algorithm performing an incomplete Gröbner basis computation which is stopped when all S-polynomials of degree $\leq d_{1}$ have been consider at a current step. This algorithm corresponds having a bound on the size of the Macaulay matrix when using linear algebra methods for computing the Gröbner basis. If the degree $d_{1}$ is sufficiently large, we have that $G={\textsc{Grobner}}(H,d_{1})$ is a complete Gröbner basis of $J+{\mathcal{F}}$. In this case, we say that $d_{1}$ is a solving degree for the ideal $J+{\mathcal{F}}$ with respect to the given monomial ordering. However, if $d_{1}<{\mathrm{maxdeg}}(H)$, then $G={\textsc{Grobner}}(H,d_{1})$ is generally not a generating set of the ideal $J+{\mathcal{F}}$. Instead, a generating set is given by $G\cup G^{\prime}$, where $G^{\prime}=\{g\in H\mid\deg(g)>d_{1}\}$.
A useful procedure within the solving algorithm we are about to propose, is GBElimLin. Its goal is to generate linear polynomials belonging to the ideal $J+{\mathcal{F}}$ in order to eliminate variables from its generators. The method consists of computing an incomplete Gröbner basis up to a sufficiently low degree $d_{1}$ to keep the computation efficient. The expectation that linear polynomials appear among the generators arises from the fact that the complete Gröbner basis is entirely linear when #$\#V_{\mathbb{F}}(J)=1$. Note also that, when incorporating this procedure into the multistep solving algorithm, we explicitly add linear polynomials corresponding to variable evaluations to the generating set.
@@
In the GBElimLin algorithm, we assume that the linear polynomials in the set $L$ are fully interreduced, meaning they are in reduced row echelon form in the Macaulay matrix used to compute ${\textsc{Grobner}}(G,d_{1})$. The Reduce procedure implements the full reduction algorithm from Gröbner basis theory. Note that if ${\mathrm{maxdeg}}(G)\leq 1$, then $G$ is a Gröbner basis of $J+{\mathcal{F}}$. In other words, GBElimLin directly returns $V_{\mathbb{F}}(J)$, as $d_{1}$ is a solving degree in this case. Finally, note that a modified version of GBElimLin could be iteratively applied to enhance its capability to eliminate variables.
To describe our solving algorithm, it is also helpful to consider a variant of Grobner with a timeout $\tau$ to limit the computation time. Specifically, we define the following procedure.
Algorithm 2 GrobnerSafe
@@
Output: either ${\tt wild},\emptyset$ or ${\tt tame},G$ where $G={\textsc{Grobner}}(H,d_{2})$ and
@@
compute $G:={\textsc{Grobner}}(H,d_{2})$ within timeout $\tau$;
@@
return ${\tt wild},\emptyset$;
@@
If $V_{\mathbb{F}}(J)=\{c_{i_{1}},\ldots,c_{i_{n}}\}$, we have that
$V_{\mathbb{F}}(J(j_{1},\ldots,j_{k}))=\left\{\begin{array}[]{cl}\{c_{i_{1}},\ldots,c_{i_{n}}\}&\mbox{if}\ (j_{1},\ldots,j_{k})=(i_{1},\ldots,i_{k});\\ \emptyset&\mbox{otherwise}.\end{array}\right.$
@@
Let $H$ be a generating set of the ideal $J+{\mathcal{F}}$ such that #$\#V_{\mathbb{F}}(J)\leq 1$. Let Oracle be a function that takes as input a set of polynomials and an integer $k$, and returns either the string tame or the string wild. Finally, fix two integers $d_{1}\geq 0$ and $d_{2}\geq{\mathrm{maxdeg}}(H)$. The set $H$, along with the function Oracle and the integers $d_{1},d_{2}$, are global variables for the following algorithm MultiSolve which applies the multistep solving strategy to the ideal $J(i_{1},\ldots,i_{k})$. For $k=0$, we have $J()=J$ by definition.
@@
     ${\tt status},G:={\textsc{GrobnerSafe}}(G,d_{2},\tau)$;
@@
Besides the DFS implementation, a key difference of the above algorithm with the one presented in [13] is the use of a general oracle function that predicts when ${\textsc{GrobnerSafe}}(G,d_{2},\tau)$ would return ${\tt status}={\tt wild}$. In the previous version of MultiSolve, given an integer $B>0$, the oracle function was simply defined as
${\textsc{OracleNRV}}(G,k)=\left\{\begin{array}[]{ll}{\tt wild}&\mbox{if the number of remaining variables in $G$ is $>B$};\\ {\tt tame}&\mbox{otherwise}.\end{array}\right.$
@@ 3. Complexity
As a result, the execution time of MultiSolve is the sum of the time spent by GrobnerSafe handling wild and tame cases when ${\textsc{Oracle}}={\tt tame}$. We refer to cases corresponding to the status returned by GrobnerSafe as computed wild and computed tame cases. Similarly, cases determined by the values of the function Oracle are referred to as predicted wild and predicted tame cases.
@@
In particular, when $q=2$, se have $N=2M-1$ and $M=L-1$.
@@
Assume that the first formula holds for any $q$-ary tree with strictly fewer than $L$ leaves. Consider a $q$-ary tree $T$ having $N$ nodes and $L$ leaves. This tree can by obtained by a smaller tree $T^{\prime}$ by converting one of its leaves into an internal node and attaching $q$ new leaves to it. Let $N^{\prime}$ and $L^{\prime}$ be the number of nodes and leaves of $T^{\prime}$, respectively. Then, we have
@@
Let Oracle be any function and assume that the depth of computed tame cases ranges from a minimum of $B$ to a maximum of $C$. The MultiSolve algorithm implementing ${\textsc{OracleH}}_{C}$, which corresponds to the standard hybrid strategy, computes exactly $q^{C}$ tame cases. Since it is a accurate oracle, we have no computed wild cases and the number of predicted wild cases is
@@
Note that in the above result, we consider the total number of wild cases, including those that are computed by GrobnerSafe, which incur some cost. If the function Oracle is accurate, such as OracleNRV in the attack on Trivium [13], there is no such cost. However, if this oracle is not applicable, for instance in the sparse case where the number of remaining variables is insufficient to accurately predict the output status of GrobnerSafe computations, different features of the input polynomial systems has to be considered to design an accurate oracle. Developing such oracle functions is generally a challenging task, which we plan to pursue in future work with the aid of Machine Learning (ML) techniques.
An oracle function that is always available is the following
@@
This constant tame function implies that all wild cases are computed ones. By definition, a perfect oracle function has the same computed tame cases as OracleT and no computed wild cases. In other words, for a fixed choice of the parameters $d_{1},d_{2}$ and $\tau$, a perfect oracle achieves optimal MultiSolve performance in terms of the number of GrobnerSafe executions, that is, computed wild and tame cases.
Note that we may also define a constant wild function, which reduces MultiSolve to the recursive application of GBElimLin. For $d_{1}=0$, this corresponds to brute-force solving by evaluating all possible values of the variables in the base finite field ${\mathbb{F}}$. By applying Theorem 3.3, one obtains the following result.
@@
Let $N$ denote the number of GrobnerSafe executions in the MultiSolve algorithm using OracleT, and let $L$ be the corresponding number when using a perfect oracle function. The resulting speed-up ratio $N/L$ is given by
$\frac{N}{L}=\frac{qL-1}{L(q-1)}$
In particular, for $q=2$, we have $N/L=2-1/L$.
@@
Recall that the recursive calls of MultiSolve define a $q$-ary tree. For the function OracleT, these calls correspond exactly to executions of GrobnerSafe, and therefore their number $N$ can be expressed as
@@
where $L$ is the number of computed tame cases. This $L$ also represents the total number of GrobnerSafe executions when using a perfect oracle. Computing the ratio $N/L$ yields the formula of the claim. ∎
@@ 4.1. Substitution Layer $\bm{\pi}$
The substitution layer introduces nonlinearity via bitwise Toffoli gates across the word-level structure. The transformation $\pi$ operates on three words at a time, following the update rule $(a,b,c)\mapsto(a,b,c\oplus a\wedge b)$. Note that we use standard Boolean notation, where $\oplus$ denotes the XOR operation, i.e., addition in ${\mathbb{F}}={\mathrm{GF}}(2)$, and $\wedge$ denotes the AND operation, corresponding to multiplication in ${\mathbb{F}}$. Applied to the full state, $\pi$ updates the words as:
@@ 6. An algebraic attack to Aradi based on the multistep strategy
The multistep strategy offers a more automated algorithm that, starting from a reasonable initial set of guessed variables, explores the algebraic structure of the polynomial system and extends this set only when strictly necessary. The effectiveness of this approach is illustrated in Section 6.
@@
Through a series of experiments on Aradi, we identified effective parameter settings for MultiSolve in the context of this attack: specifically, $d_{1}=0$, which disables GBElimLin, and $d_{2}=2$ for GrobnerSafe. This choice is motivated by the fact that the input polynomials are quadratic, which enables GrobnerSafe to be applied efficiently at this degree while still producing complete Gröbner bases, thus rendering GBElimLin unnecessary.
@@
We report here the average runtimes of GrobnerSafe in solving tame cases, along with the corresponding standard deviations to highlight the stability of the results. All timings are given in seconds. All computations were performed on a Linux-based system running Ubuntu 22.04 LTS, equipped with an AMD EPYC 7763 64-core processor and x86_64 architecture. For Gröbner basis computations, we used the sparse F4 algorithm as implemented in Magma version 2.27.5.
@@
Since $\log_{2}(34.62)=5.11$ and $\log_{2}(24.95)=4.64$, the overall complexity of the multistep strategy is approximately $2^{256}$ seconds, while that of the classical hybrid strategy is around $2^{256.5}$ seconds. Given that a single encryption in a brute force attack takes significantly less than one second, we cannot claim that Aradi is broken. However, the effective key security is certainly subject to concern.
@@
We also experimented with various combinations of the parameters $d_{1},d_{2}$ and $\tau$, but none of these configurations enabled us to obtain tame cases with only 250 key variable evaluations. In other words, we identified a structural weakness in the algebraic representation of Aradi, manifesting as a sharp complexity drop once 251 key variables are fixed.

### v3

@@ Abstract.
In this paper, we present a new recursive formulation of the corresponding algorithm based on a Depth-First Search strategy, along with a novel complexity analysis leveraging tree structures. We also introduce the notion of an “oracle function”, which is intended to determine whether evaluating a new variable is required to simplify the current polynomial system. This notion allows us to unify all previously proposed variants of the multistep strategy, including the classical hybrid approach, by appropriately selecting the oracle function.
Finally, we employ the multistep solving strategy in the cryptanalysis of the NSA’s recently introduced low-latency block cipher Aradi, achieving a first full-round algebraic attack that exposes structural features in its symbolic model.
@@
Primary 13P15. Secondary 11T71, 11T06
The first author acknowledges the partial support of PNRR MUR projects “Security and Rights in the CyberSpace”, Grant ref. CUP H93C22000620001, Code PE00000014, Spoke 3, and “National Center for HPC, Big Data and Quantum Computing”, Grant ref. CUP H93C22000450007, Code CN00000013, Spoke 10. The same author was co-funded by MUR project PRIN 2022SC, Grant ref. 2022RFAZCJ, and Università di Bari “Fondo acquisto e manutenzione attrezzature per la ricerca”, Grant ref. DR 3191.
@@ 1. Introduction
Beyond theoretical worst-case complexity, valid when the polynomials in the system are randomly chosen, it is especially important in cryptanalysis to estimate the actual computational cost of solving a specific polynomial system over a finite field through experimentation. Since solving such systems becomes infeasible when the number of variables is sufficiently large, a standard approach is to combine brute force with polynomial solving in what is known as a “hybrid strategy”. Specifically, this approach assumes that one can feasibly solve all derived systems obtained by exhaustively assigning values in the finite field to $k$ out of the $n$ variables in the original system. By using a test set of such partial evaluations, it is possible to estimate the average solving time $\tau$ for a single derived system. Consequently, the total expected complexity of the hybrid approach becomes $q^{k}\tau$.
@@
Moreover, we observe that all derived systems are quadratic and can be solved with a solving degree of two when more than 250 key variable assignments are provided. However, this is no longer possible when the number of assignments is reduced to 250. These findings point to some thresholds within Aradi’s algebraic structure, and they show that a dynamic multistep strategy can be practically useful in algebraic cryptanalysis.
@@ 2. Multistep algorithm
Let ${\mathbb{F}}={\mathrm{GF}}(q)$ be a finite field and consider the multivariate polynomial algebra $R={\mathbb{F}}[x_{1},\ldots,x_{n}]$ and the field equation ideal ${\mathcal{F}}=\langle x_{1}^{q}-x_{1},\ldots,x_{n}^{q}-x_{n}\rangle\subset R$. Let $J=\langle f_{1},\ldots,f_{m}\rangle\subset R$ be an ideal such that $V_{\mathbb{F}}(J)\leq 1$. Put
@@
that is, $H$ is a generating set of the ideal $J+{\mathcal{F}}$. To simplify notations, we denote ${\mathbb{F}}=\{c_{1},\ldots,c_{q}\}$. Since $V_{\mathbb{F}}(J)\leq 1$, we have that a (reduced) Gröbner basis of $J+{\mathcal{F}}$ with respect to any monomial ordering of $R$, is either $G=\{1\}$ or $G=\{x_{1}-c_{i_{1}},\ldots,x_{n}-c_{i_{n}}\}$, for some $1\leq i_{1},\ldots,i_{n}\leq q$ (see, for instance, [10, 13, 15]). In other words, ${\mathrm{maxdeg}}(G)=\max\{\deg(f)\mid f\in G\}\leq 1$.
Let $d_{1}\geq 0$ be an integer and denote by ${\textsc{Groebner}}(H,d_{1})$ the algorithm performing an incomplete Gröbner basis computation which is stopped when all S-polynomials of degree $\leq d_{1}$ have been consider at a current step. This algorithm corresponds having a bound on the size of the Macaulay matrix when using linear algebra methods for computing the Gröbner basis. If the degree $d_{1}$ is sufficiently large, we have that $G={\textsc{Groebner}}(H,d_{1})$ is a complete Gröbner basis of $J+{\mathcal{F}}$. In this case, we say that $d_{1}$ is a solving degree for the ideal $J+{\mathcal{F}}$ with respect to the given monomial ordering. However, if $d_{1}<{\mathrm{maxdeg}}(H)$, then $G={\textsc{Groebner}}(H,d_{1})$ is generally not a generating set of the ideal $J+{\mathcal{F}}$. Instead, a generating set is given by $G\cup G^{\prime}$, where $G^{\prime}=\{g\in H\mid\deg(g)>d_{1}\}$.
A useful procedure within the solving algorithm is GBElimLin. Its goal is to generate linear polynomials belonging to the ideal $J+{\mathcal{F}}$ in order to eliminate variables from its generators. The method consists of computing an incomplete Gröbner basis up to a sufficiently low degree $d_{1}$ to keep the computation efficient. The expectation that linear polynomials appear among the generators arises from the fact that the complete Gröbner basis is entirely linear when $V_{\mathbb{F}}(J)=1$. Note also that, when incorporating this procedure into the multistep solving algorithm, we explicitly add linear polynomials corresponding to variable evaluations to the generating set.
@@
In the GBElimLin algorithm, we assume that the linear polynomials in the set $L$ are fully interreduced, meaning they are in reduced row echelon form in the Macaulay matrix used to compute ${\textsc{Groebner}}(G,d_{1})$. The Reduce procedure implements the full reduction algorithm from Gröbner basis theory. Note that if ${\mathrm{maxdeg}}(G)\leq 1$, then $G$ is a Gröbner basis of $J+{\mathcal{F}}$. In other words, GBElimLin directly returns $V_{\mathbb{F}}(J)$, as $d_{1}$ is a solving degree in this case. Finally, note that a modified version of GBElimLin could be iteratively applied to enhance its capability to eliminate variables.
To describe our solving algorithm, it is also helpful to consider a variant of Groebner with a timeout $\tau$ to limit the computation time. Specifically, we define the following procedure.
Algorithm 2 GroebnerSafe
@@
Output: either $\tt wild$$\emptyset$ or ${\tt tame},G$ where $G={\textsc{Groebner}}(H,d_{2})$ and
@@
compute $G:={\textsc{Groebner}}(H,d_{2})$ within timeout $\tau$;
@@
return $\tt wild$$\emptyset$;
@@
If $V_{\mathbb{F}}(J)=\{(c_{i_{1}},\ldots,c_{i_{n}})\}$, we have that
$V_{\mathbb{F}}(J(j_{1},\ldots,j_{k}))=\left\{\begin{array}[]{cl}V_{\mathbb{F}}(J)&\mbox{if}\ (j_{1},\ldots,j_{k})=(i_{1},\ldots,i_{k});\\ \emptyset&\mbox{otherwise}.\end{array}\right$.
@@
Let $H$ be a generating set of the ideal $J+{\mathcal{F}}$ such that $V_{\mathbb{F}}(J)\leq 1$. Let Oracle be a function that takes as input a set of polynomials and an integer $k$, and returns either the string tame or the string wild. Finally, fix two integers $d_{1}\geq 0$ and $d_{2}\geq{\mathrm{maxdeg}}(H)$. The set $H$, along with the function Oracle and the integers $d_{1},d_{2}$, are global variables for the following algorithm MultiSolve which applies the multistep solving strategy to the ideal $J(i_{1},\ldots,i_{k})$. For $k=0$, we have $J()=J$ by definition.
@@
  ${\tt status},G:={\textsc{GroebnerSafe}}(G,d_{2},\tau)$;
@@
Besides the DFS implementation, a key difference of the above algorithm from the one presented in [14] is the use of a general oracle function meant to predict when ${\textsc{GroebnerSafe}}(G,d_{2},\tau)$ would return ${\tt status}={\tt wild}$. In the previous version of MultiSolve, given an integer $B>0$, the oracle function was simply defined as
[…]
@@
We remark that the MultiSolve algorithm essentially imposes a fixed solving degree $d_{2}$, at the cost of performing branching among the various derived polynomial systems. Combined with its recursive structure, this provides MultiSolve with considerable simplicity and flexibility, which can be adjusted by varying the parameters $d_{1},d_{2},\tau$ and the function Oracle.
@@ 3. Complexity
As a result, the execution time of MultiSolve is the sum of the time spent by GroebnerSafe handling wild and tame cases when ${\textsc{Oracle}}={\tt tame}$. We refer to cases corresponding to the status returned by GroebnerSafe as computed wild and computed tame cases. Similarly, cases determined by the values of the function Oracle are referred to as predicted wild and predicted tame cases.
@@
In particular, when $q=2$, se have N=2L1N=2L-1 and M=L1M=L-1.
@@
Assume that the first formula holds for any $q$-ary tree with strictly fewer than $L$ leaves. Consider a $q$-ary tree $T$ having $N$ nodes and $L$ leaves. This tree can be obtained by a smaller tree $T^{\prime}$ by converting one of its leaves into an internal node and attaching $q$ new leaves to it. Let $N^{\prime}$ and $L^{\prime}$ be the number of nodes and leaves of $T^{\prime}$, respectively. Then, we have
@@
Let Oracle be any function and assume that the depth of computed tame cases ranges from a minimum of $B$ to a maximum of $C$. The MultiSolve algorithm implementing ${\textsc{OracleH}}_{C}$, which corresponds to the standard hybrid strategy, computes exactly $q^{C}$ tame cases. Since it is an accurate oracle, we have no computed wild cases and the number of predicted wild cases is
@@
Note that in the above result, we consider the total number of wild cases, including those that are computed by GroebnerSafe, which incur some cost. If the function Oracle is accurate, such as OracleNRV in the attack on Trivium [14], there is no such cost. However, if this oracle is not applicable, for instance in the sparse case where the number of remaining variables is insufficient to accurately predict the output status of GroebnerSafe computations, different features of the input polynomial systems has to be considered to design an accurate oracle. Developing such oracle functions is generally a challenging task, which we plan to pursue in future work with the aid of Machine Learning (ML) techniques.
The following is a trivial, non-predictive oracle that is always available
@@
Indeed, this constant tame function implies that all wild cases are computed ones. By definition, a perfect oracle function has the same computed tame cases as OracleT and no computed wild cases. In other words, for a fixed choice of the parameters $d_{1},d_{2}$ and $\tau$, a perfect oracle achieves optimal MultiSolve performance in terms of the number of GroebnerSafe executions, that is, computed wild and tame cases. At the opposite end of the spectrum, the function OracleT yields the worst performance.
By applying Theorem 3.3, one obtains the following result.
@@
Let $N$ denote the number of GroebnerSafe executions in the MultiSolve algorithm using OracleT, and let $L$ be the corresponding number when using a perfect oracle function. The resulting speed-up ratio $N/L$ is given by
$\frac{N}{L}=\frac{qL-1}{L(q-1)}=1+\frac{1}{q-1}(1-\frac{1}{L})$
In particular, for $q=2$, we have N/L=21/LN/L=2-1/L, which implies $1\leq N/L<2$.
@@
Recall that the recursive calls of MultiSolve define a $q$-ary tree. For the function OracleT, these calls correspond exactly to executions of GroebnerSafe, and therefore their number $N$ can be expressed as
@@
where $L$ is the number of computed tame cases. This $L$ also represents the total number of GroebnerSafe executions when using a perfect oracle. Computing the ratio $N/L$ yields the formula of the claim. ∎
Note that we may also define a constant wild function, which reduces MultiSolve to the recursive application of GBElimLin. For $d_{1}=0$, this corresponds to brute-force solving by evaluating all possible values of the variables in the base finite field $\mathbb{F}$}.
@@ 4.1. Substitution Layer $\bm{\pi}$
The substitution layer introduces nonlinearity via bitwise Toffoli gates across the word-level structure. The transformation $\pi$ operates on three words at a time, following the update rule $(a,b,c)\mapsto(a,b,c\oplus(a\wedge b))$. Note that we use standard Boolean notation, where $\oplus$ denotes the XOR operation, i.e., addition in ${\mathbb{F}}={\mathrm{GF}}(2)$, and $\wedge$ denotes the AND operation, corresponding to multiplication in $\mathbb{F}$}. Applied to the full state, $\pi$ updates the words as:
@@ 6. An algebraic attack to Aradi based on the multistep strategy
The multistep strategy offers a more automated algorithm that, starting from a reasonable initial set of guessed variables, explores the algebraic structure of the polynomial system and extends this set only when strictly necessary.
@@
Through a series of experiments on Aradi, we identified effective parameter settings for MultiSolve in the context of this attack: specifically, $d_{1}=0$, which disables GBElimLin, and $d_{2}=2$ for GroebnerSafe. This choice is motivated by the fact that the input polynomials are quadratic, which enables GroebnerSafe to be applied efficiently at this degree while still producing complete Gröbner bases, thus rendering GBElimLin unnecessary.
@@
We report here the average runtimes of GroebnerSafe in solving tame cases, along with the corresponding standard deviations to highlight the stability of the results. All timings are given in seconds. All computations were performed on a Linux-based system running Ubuntu 22.04 LTS, equipped with an AMD EPYC 7763 64-core processor and x86_64 architecture. For Gröbner basis computations, we used the sparse F4 algorithm as implemented in Magma version 2.27.5.
@@
Since $\log_{2}(34.62)=5.11$ and $\log_{2}(24.95)=4.64$, the overall complexity of the multistep strategy is approximately $2^{256}$ seconds, while that of the classical hybrid strategy is around $2^{256.5}$ seconds. Given that a single encryption in a brute force attack takes significantly less than one second, we cannot claim that Aradi is broken. However, our attack has revealed some structural features of the symbolic model.
@@
We also experimented with various combinations of the parameters $d_{1},d_{2}$ and $\tau$, but none of these configurations enabled us to obtain tame cases with only 250 key variable evaluations. In other words, fixing 251 key variables in Aradi’s symbolic model results in a marked drop in computational effort.